There are 12 teams in a curling competition. After the round-robin portion the 1st place teams plays the 4th place team and the 2nd place team plays the 3rd place team in the semi-finals. Determine the number of different possible pairings for the semifinal games. Explain your answer.
Any help would be appreciated. Thanks.

Answer by stanbon(75887)

There are 12 teams in a curling competition. After the round-robin portion the 1st place teams plays the 4th place team and the 2nd place team plays the 3rd place team in the semi-finals. Determine the number of different possible pairings for the semifinal games. Explain your answer.
--------------------------
Choose two teams to be the 1st and 4th in 12C2 = 12*11/1*2 = 66 ways
--------------------------------
Choose two teams to be the 2nd and 3rd in 10C2 = 10*9/1/2 = 45 ways
--------------------------------
Total number of pairings for the semifinals = 66*45 = 2970
===============
Cheers,
Stan H.